OpenStudy (anonymous):

solve the question x^2+18-10

OpenStudy (anonymous):

\[\frac{ -b \frac{ + }{ }\sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

There's nothing to solve

OpenStudy (anonymous):

sorry i meant by competing the square

OpenStudy (anonymous):

omg i thought there was an x on the 18 sorry

OpenStudy (anonymous):

that too 18x not 18

OpenStudy (anonymous):



OpenStudy (anonymous):

o ok